# Towards Constraint-aware Learning for Resource Allocation in NFV-enabled Networks

### v1

@@ B.1 Optimization Objectives
The main goal of VNE is to make full use of the physical network resources to improve the revenue of ISPs while satisfying the service requests of users as much as possible. To address the stochastic nature of online networking, we and existing studies (Wang et al., 2021; He et al., 2023a; Zhang et al., 2023b), aim to minimize the embedding cost of each arriving VN request onto the physical network. This way enhances resource utilization and improves the VN request acceptance rate. To evaluate the quality of solution $E=f_{G}(I)$, we employ the widely used indicator, Revenue-to-Consumption ratio (R2C), defined as follows:
@@ B.2 Constraint Conditions
Here, $\Omega(n_{p}^{k})$ denotes the neighbors of $n_{p}^{k}$. Constraint (10) ensures that every virtual node is mapped to one and only one physical node. Conversely, constraint (11 )limits each physical node to hosting at most one virtual node, thus enforcing a unique mapping relationship. Constraint (12) verifies that virtual nodes are allocated to physical nodes with adequate resources. Following the principle of flow conservation, constraint 13 guarantees that each virtual link $(n_{v}^{m},n_{v}^{w})$ is routed along a physical path from $n_{p}^{i}$ (the physical node where $n_{v}^{m}$ is placed) to $n_{p}^{j}$ (the physical node where $n_{v}^{w}$ is placed). Constraint (14) eliminates the possibility of routing loops, thereby ensuring that virtual links are routed acyclically. Lastly, constraint 15 ensures that the bandwidth usage on each physical link remains within its available capacity. Overall, constraints (10,11,12) enforce the one-to-one placement and computing resource availability required in the node mapping process. And constraints (13,14,15) ensure the path connectivity and bandwidth resource availability asked in the link mapping process.
@@ Appendix C Preliminary Study
We have conducted a preliminary study to highlight the motivation to handle failure samples and unsolvable instances. In this study, we trained a representative baseline model, A3C-GCN Zhang et al. (2023b), which shares a solution construction paradigm similar to our approach and uses a penalty mechanism to handle all failure samples. The training was conducted with various arrival rates for VN requests because it is evident that increasing the arrival rate of VN requests leads to a higher frequency of unsolvable instances. After training, we tested the models using a fixed arrival rate $\lambda=0.14$ (the same as in Section 4.1: Experimental Settings) with a seed of 0 as the benchmark.
@@ D.2 Proof of Lagrange Multiplier Convergence
$H(s)$ $\leq 0,$ (16)
$\lambda$ $\geq 0,$
$\lambda\cdot H(s)$ $=0.$
Given that $H(s)>0,\forall s\in S$, the primal feasibility condition is violated; there are no feasible solutions that satisfy $H(s)\leq 0$. The complementary slackness condition requires that for each $s\in S$, either $\lambda=0$ or $H(s)=0$. Since $H(s)>0$, we must have $\lambda=0$ to satisfy this condition. However, setting $\lambda=0$ does not penalize the constraint violation, and the primal feasibility condition remains unsatisfied. This leads to a contradiction: there is no finite $\lambda\geq 0$ that satisfies all the KKT conditions when $H(s)>0,\forall s\in S$. The optimization problem is infeasible because the constraints cannot be met. In practical online training, the Lagrange multiplier $\lambda$ is adjusted iteratively to enforce the constraints by increasing $\lambda$ whenever the constraints are violated. Since $H(s)>0$ always holds, the Lagrange multiplier $\lambda$ will continually increase in an attempt to penalize the constraint violations. As a result, $\lambda$ can become arbitrarily large, theoretically approaching infinity. ∎
@@ D.4 Barlow Twins Loss Function
In this work, we utilize the contrastive learning method to bring the representations under augmented views close to enhance the awareness of bandwidth-constrained path connectivity. One of the main directions of contrastive learning is to utilize both positive and negative sample distinctions (Oord et al., 2018; Chen et al., 2020). However, in our application, generating and selecting negative samples that are markedly different in terms of feasibility semantics is significantly challenging. This difficulty can adversely affect the learning quality and the generalizability of the models. Therefore, considering multiple views of the graph with the same feasibility semantics, we focus on eschewing negative samples altogether while preventing feature collapse, an emerging direction of contrastive learning (Grill et al., 2020; Zbontar et al., 2021). Specifically, we adopt the Barlow Twins method (Zbontar et al., 2021) for its simplicity and effectiveness, which circumvents negative sample selection and maintains the original network architecture. Given the node embeddings under two augmented views, $H^{a}$ and $H^{b}$, we reduce redundancy between embedding components by aligning their cross-correlation matrix with the identity matrix:
$L_{\textit{CL}}=\sum_{i}\left(1-\mathbb{C}_{ii}\right)^{2}+w\sum_{i}\sum_{i\neq j}\mathbb{C}_{ij}^{2},$ (23)
@@
$L_{\textit{LAM}}=\Lambda^{\pi}(s)\cdot(V_{h}^{\pi}(s)-D^{\pi^{\prime}}_{h}(s)).$ (24)
@@
$L_{\textit{PPO}}(\pi)=\mathbb{E}[\min(r(\pi)A_{t},\text{CLIP}(r(\pi),1-\epsilon,1+\epsilon)A_{t})],$ (25)
where $A_{t}$ represents the advantage function at time step $t$, and $r(\pi)=\frac{\pi(s_{t},a_{t})}{\pi{{}^{\textit{old}}}(s_{t},a_{t})}$ calculates the ratio of the probabilities for the selected action between the current and previous policies. The CLIP function limits policy updates to enhance stability.
Finally, integrating the unsupervised contrastive learning objective $L_{\textit{CL}}$ in the path-bandwidth contrast method, our comprehensive loss function in the training process is formulated as follows:
$L=w_{\textit{PPO}}\cdot L_{\textit{PPO}}+w_{\textit{r}}\cdot L_{r}+w_{h}\cdot L_{h}+w_{\textit{LAM}}\cdot L_{\textit{LAM}}+w_{\textit{CL}}\cdot L_{\textit{CL}},$ (26)
@@ D.6 Detailed Analysis of Computational Complexity
CONAL exhibits a computational complexity of $O\left(|N_{v}|\cdot K\cdot\left(|L_{p}|d+|N_{p}+N_{v}|d^{2}\right)\right)$, which the complexities other baseline methods based on RL and GNNs are $O\left(|N_{v}|\cdot K\cdot\left(|L_{p}|d+|N_{p}|d^{2}\right)\right)$. Here, $N_{v}$ and $L_{v}$ denote the number of virtual nodes and links, $N_{p}$ and $L_{p}$ denote the number of physical nodes and links, $K$ denotes the number of GNN layers, and $d$ denotes the embedding dimension. Concretely, When constructing a solution for one VNE instance, CONAL performs inference $N_{v}$ times with the GNN policy, similar to most RL and GNN-based methods. The difference in complexity between CONAL and RL/GNN-based baselines mainly arises from the different neural network structures used, such as GAT and GCN. One GAT and one GCN layer have the same complexity, both $O\left(|L|d+|N|d^{2}\right)$, where $|N|$ and $|L|$ denote the number of nodes and links [a]. In CONAL, we enhance the GAT with the heterogeneous modeling for the interactions of cross-graph status. Each heterogeneous GAT layer consists of three types of GAT layers for VN, PN, and cross-graph interactions (the number of links between virtual and physical nodes is always $N_{p}$). The complexities for these layers are $O\left(|L_{v}|d+|N_{v}|d^{2}\right)$, $O\left(|L_{p}|d+|N_{p}|d^{2}\right)$, and $O\left(|N_{p}|d+|N_{p}+N_{v}|d^{2}\right)$, respectively. Each heterogeneous GAT layer consists of three types of GAT layers for VN, PN, and cross graph (the number of links between virtual and physical node always is $N^{p}$), whose complexity is $O\left(|L_{v}|d+|N_{v}|d^{2}\right)$, $O\left(|L_{p}|d+|N_{p}|d^{2}\right)$, and $O\left(|N_{p}|d+|N_{p}+N_{v}|d^{2}\right)$. Considering that $|N_{v}|$ is significantly smaller than $|L_{v}|$ and typically $L_{p}>N_{p}$ in practical network systems, the overall complexity of CONAL is $O\left(|N_{v}|\cdot K\cdot\left(|L_{p}|d+|N_{p}+N_{v}|d^{2}\right)\right)$. In comparison, other RL and GNN-based methods separately encode VN and PN with GAT or GCNs, without considering GNN layers for cross-graph interactions, leading to their complexities being $O\left(|N_{v}|\cdot K\cdot\left(|L_{p}|d+|N_{p}|d^{2}\right)\right)$. Overall, while CONAL slightly increases the complexity compared to existing RL and GNN-based methods due to its heterogeneous modeling approach, it achieves significant performance improvements.
@@ Appendix E Descriptions of Training and Inference Process
In this section, we summarize the training process of CONAL in Algorithm 1 and the inference process of CONAL in Algorithm 2.
@@ E.1 Training Process of CONAL
19:     Calculate the PPO objective $L_{\textit{PPO}}$ for policy $\pi$ with Eq. 25
@@
22:     Calculate the contrastive loss $L_{\textit{CL}}$ with the Barlow Twins method, i.e., Eq. 23
@@
In the subsequent policy optimization stage, we iteratively sample batches of transitions from the trajectory memory $\tau$ to update neural networks. Then, we calculate various losses, including the PPO loss $L_{\textit{PPO}}$, value critic loss $L_{r}$, reachability critic loss $L_{h}$, and Lambda network loss $L_{\textit{LAM}}$. Additionally, we integrate our proposed unsupervised contrastive loss $L_{\textit{CL}}$ in path-bandwidth contrast during training. Overall, we calculate the weighted sum of these losses as the final loss $L$ and update neural networks. If necessary, we synchronize surrogate policy $\pi^{\prime}$ with the main policy $\pi$.
@@ E.2 Inference Process of CONAL
During the inference process, we use the learned policy $\pi$ to solve newly arrived VNE instances. Note that the surrogate policy and the path-bandwidth contrast module are not utilized in this process. The inference process is outlined in Algorithm 2.
@@ F.1 Implementation Details
Simulation for Training and Testing. For RL-based methods and CONAL, we train policies for each average arrival rate $\eta$, where running seeds are randomly set in every simulation. During testing, we evaluate the performance of all algorithms by repeating the tests with 10 different seeds (i.e., $0,1111,2222,\cdots,9999$) for each average arrival rate $\eta$ to ensure statistical significance.
@@ F.2 Baseline Descriptions
  • PG-CNN (Zhang et al., 2022) is a model-free RL method. It models the solution construction of each VNE instance as MDP. Then, it develops a policy network with Convolutional Neural Network (CNN) and trains it using the Policy Gradient (PG) algorithm. Specifically, during training, only samples related to feasible solutions are used for optimization.
  • A3C-GCN (Zhang et al., 2023b) is a model-free RL method. It constructs a policy network with a Graph Convolutional Network (GCN) and a Multi-Layer Perceptron (MLP). The Asynchronous Advantage Actor-Critic (A3C) algorithm is used for training. Particularly, during training, if encountering failure samples, customized negative rewards are returned.
  • DDPG-ATT (He et al., 2023a) is a model-free RL method that builds an ATTention-based (ATT) policy network and uses the Deep Deterministic Policy Gradient (DDPG) algorithm for training. Similar to A3C-GCN, it introduces the penalty into reward shaping.
  • GAL-VNE (Geng et al., 2023) is a model-free RL method that formulates VNE into both Global learning across requests And Local prediction within requests (GAL). It first uses supervised learning to develop a GNN policy to solve VNE in one shot. Then, in online perspective, they use the RL to finetune this policy to improve the overall performance.
@@ F.3 Variations Descriptions
  • $\text{CONAL}_{\text{w/o ARB}}$ removes the proposed adaptive reachability budget (ARB) module and uses zero as a fixed reachability budget.
@@ Appendix G Addtional Evaluation
In this section, we present additional experiments to evaluate the generalizability, scalability, and practicability of CONAL, as well as investigate the impact of two key hyperparameters.
G.1 Generalizability Study
@@
Dynamic Request Distribution Testing. In practical scenarios, the node size and resource requirements distributions of VN requests may vary due to different service demands. To simulate such situations, similar to previous work (Geng et al., 2023), we equally divided 1000 VN requests into four sub-groups. In comparison to the default simulation settings, we modified one parameter related to the distribution of VN resource demand or VN node size for each sub-group: For the first group, the node and link resource distributions of VN are changed to [0, 30] and [0, 75], respectively; For the second group, the node and link resource distributions of VN are adjusted to [0, 40] and [0, 100], respectively; For the third group, the VN node size distribution is changed to [2, 15]; For the fourth group, the VN node size distribution is altered to [2, 20]. We evaluate the CONAL and other RL-based methods trained in the default settings. The results are illustrated in Figure 6. We observe that CONAL exhibits superior performance across all metrics, indicating its strong adaptability to dynamic request distributions. Regarding the VN_ACR, all algorithms experience a rapid decrease in the early stages of the simulation due to the quick consumption of resources. Subsequently, their VN_ACR becomes relatively stable. It is worth noting that in the fourth stage, most algorithms show a clear downward trend in VN_ACR. This is because embedding larger VNs with more complex topologies and greater resource demands becomes increasingly challenging. This testing shows that CONAL is highly adaptable across varying VN request distributions, which underscores CONAL’s generalization in dynamic network environments.

### v2

@@ B.1 Optimization Objectives
The main goal of VNE is to make full use of the physical network resources to improve the revenue of ISPs while satisfying the service requests of users as much as possible. To address the stochastic nature of online networking, we and existing studies (Wang et al., 2021; He et al., 2023a; Zhang et al., 2024b), aim to minimize the embedding cost of each arriving VN request onto the physical network. This way enhances resource utilization and improves the VN request acceptance rate. To evaluate the quality of solution $E=f_{G}(I)$, we employ the widely used indicator, Revenue-to-Consumption ratio (R2C), defined as follows:
@@ B.2 Constraint Conditions
Here, $\Omega(n_{p}^{k})$ denotes the neighbors of $n_{p}^{k}$. Constraint (10) ensures that every virtual node is mapped to one and only one physical node. Conversely, constraint (11 )limits each physical node to hosting at most one virtual node, thus enforcing a unique mapping relationship due to isolation requirements. Constraint (12) verifies that virtual nodes are allocated to physical nodes with adequate resources. Following the principle of flow conservation, constraint 13 guarantees that each virtual link $(n_{v}^{m},n_{v}^{w})$ is routed along a physical path from $n_{p}^{i}$ (the physical node where $n_{v}^{m}$ is placed) to $n_{p}^{j}$ (the physical node where $n_{v}^{w}$ is placed). Constraint (14) eliminates the possibility of routing loops, thereby ensuring that virtual links are routed acyclically. Lastly, constraint 15 ensures that the bandwidth usage on each physical link remains within its available capacity. Overall, constraints (10,11,12) enforce the one-to-one placement and computing resource availability required in the node mapping process. And constraints (13,14,15) ensure the path connectivity and bandwidth resource availability asked in the link mapping process.
@@ Appendix C Preliminary Study
We have conducted a preliminary study to highlight the motivation to handle failure samples and unsolvable instances. In this study, we trained a representative baseline model, A3C-GCN (Zhang et al., 2024b), which shares a solution construction paradigm similar to our approach and uses a penalty mechanism to handle all failure samples. The training was conducted with various arrival rates for VN requests because it is evident that increasing the arrival rate of VN requests leads to a higher frequency of unsolvable instances. After training, we tested the models using a fixed arrival rate $\lambda=0.14$ (the same as in Section 4.1: Experimental Settings) with a seed of 0 as the benchmark.
@@ D.2 Proof of Lagrange Multiplier Convergence
$H(s)$ 0$\leq$ 0,≤ 0 , (16)
$\lambda$ 0$\geq$ 0,≥ 0 ,
$\lambda\cdot H(s)$ =0=0.= 0 .
Given that $H(s)>0,\forall s\in S$, the primal feasibility condition is violated; there are no feasible solutions that satisfy $H(s)\leq 0$. The complementary slackness condition requires that for each $s\in S$, either $\lambda=0$ or $H(s)=0$. Since $H(s)>0$, we must have $\lambda=0$ to satisfy this condition. However, setting $\lambda=0$ does not penalize the constraint violation, and the primal feasibility condition remains unsatisfied. This leads to a contradiction: there is no finite $\lambda\geq 0$ that satisfies all the KKT conditions when $H(s)>0,\forall s\in S$. The optimization problem is infeasible because the constraints cannot be met. In practical online training, the Lagrange multiplier $\lambda$ is adjusted iteratively to enforce the constraints by increasing $\lambda$ whenever the constraints are violated. Since $H(s)>0$ always holds, the Lagrange multiplier $\lambda$ will continually increase in an attempt to penalize constraint violations. As a result, $\lambda$ can become arbitrarily large, theoretically approaching infinity.
@@ D.4 Barlow Twins Loss Function
In this work, we utilize the contrastive learning method to bring the representations under augmented views close to enhance the awareness of bandwidth-constrained path connectivity. One of the main directions of contrastive learning is to utilize both positive and negative sample distinctions (Oord et al., 2018; Chen et al., 2020). However, in our application, generating and selecting negative samples that are markedly different in terms of feasibility semantics is significantly challenging. This difficulty can adversely affect the learning quality and the generalizability of the models. Therefore, considering multiple views of the graph with the same feasibility semantics, we focus on eschewing negative samples altogether while preventing feature collapse, an emerging direction of contrastive learning (Grill et al., 2020; Zbontar et al., 2021). Specifically, we adopt the Barlow Twins method (Zbontar et al., 2021) for its simplicity and effectiveness, which circumvents negative sample selection and maintains the original network architecture. Given the node embeddings under two augmented views, $H^{a}$ and $H^{b}$, we reduce redundancy between embedding components by aligning their cross-correlation matrix with the identity matrix/ Specifically, we engage the subsequent optimization with following unsupervised learning loss objective function:
$L_{\textit{\tiny CL}}=\sum_{i}\left(1-\mathbb{C}_{ii}\right)^{2}+w\sum_{i}\sum_{i\neq j}\mathbb{C}_{ij}^{2},$ (23)
@@
$L_{\textit{\tiny LAM}}=\Lambda^{\pi}(s)\cdot(V_{h}^{\pi}(s)-D^{\pi^{\prime}}_{h}(s)).$ (24)
@@
$L_{\textit{\tiny PPO}}(\pi)=\mathbb{E}[\min(r(\pi)A_{t},\text{CLIP}(r(\pi),1-\epsilon,1+\epsilon)A_{t})],$ (25)
where $A_{t}$ represents the advantage function at time step $t$, and $r(\pi)=\frac{\pi(s_{t},a_{t})}{\pi{{}^{\textit{\tiny old}}}(s_{t},a_{t})}$ calculates the ratio of the probabilities for the selected action between the current and previous policies. The CLIP function limits policy updates to enhance stability.
Finally, integrating the unsupervised contrastive learning objective $L_{\textit{\tiny CL}}$ in the path-bandwidth contrast method, our comprehensive loss function in the training process is formulated as follows:
$L=w_{\textit{\tiny PPO}}\cdot L_{\textit{\tiny PPO}}+w_{\textit{\tiny r}}\cdot L_{r}+w_{h}\cdot L_{h}+w_{\textit{\tiny LAM}}\cdot L_{\textit{\tiny LAM}}+w_{\textit{\tiny CL}}\cdot L_{\textit{\tiny CL}},$ (26)
@@ D.6 Detailed Analysis of Computational Complexity
[…]
@@ Appendix E Descriptions of Training and Inference Process
In this section, we summarize the training and inference process of CONAL in Algorithm 1 and Algorithm 2, respectively.
@@ E.1 Training Process of CONAL
19:  Calculate the PPO objective $L_{\textit{\tiny PPO}}$ for policy $\pi$ with Eq. 25
@@
22:  Calculate the contrastive loss $L_{\textit{\tiny CL}}$ with the Barlow Twins method, i.e., Eq. 23
@@
In the subsequent policy optimization stage, we iteratively sample batches of transitions from the trajectory memory $\tau$ to update neural networks. Then, we calculate various losses, including the PPO loss $L_{\textit{\tiny PPO}}$, value critic loss $L_{r}$, reachability critic loss $L_{h}$, and Lambda network loss $L_{\textit{\tiny LAM}}$. Additionally, we integrate our proposed unsupervised contrastive loss $L_{\textit{\tiny CL}}$ in path-bandwidth contrast during training. Overall, we calculate the weighted sum of these losses as the final loss $L$ and update neural networks. If necessary, we synchronize surrogate policy $\pi^{\prime}$ with the main policy $\pi$.
@@ E.2 Inference Process of CONAL
Note that the surrogate policy and the path-bandwidth contrast module are not utilized in this process. During the inference process, we use the learned policy $\pi$ to solve newly arrived VNE instances, which is outlined in Algorithm 2.
@@ F.1 Implementation Details
Simulation for Training and Testing. For RL-based methods and CONAL, we train policies for each average arrival rate $\eta$, where running seeds are randomly set in every simulation. During testing, we evaluate the performance of all algorithms by repeating the tests with 10 different seeds (i.e., 0,$1$,$2$,,99990,$1$,$2$$\cdots$,99990 , $1$ , $2$ , ⋯ , $9$) for each $\eta$ to ensure statistical significance.
@@ F.2 Baseline Descriptions
  • PG-CNN (Zhang et al., 2022) is a model-free RL method. It models the solution construction of each VNE instance as MDP. Then, it develops a policy network with Convolutional Neural Network (CNN) and trains it using the Policy Gradient (PG) algorithm. To avoid unnecessary failure during action selection, a mask vector is applied to ensure that only physical nodes with sufficient computing resources are selected. Specifically, during training, only samples related to feasible solutions are used for optimization.
  • DDPG-ATT (He et al., 2023a) is a model-free RL method that builds an ATTention-based (ATT) policy network and uses the Deep Deterministic Policy Gradient (DDPG) algorithm for training.
  • A3C-GCN (Zhang et al., 2024b) is a model-free RL method. It constructs a policy network with a Graph Convolutional Network (GCN) and a Multi-Layer Perceptron (MLP). The Asynchronous Advantage Actor-Critic (A3C) algorithm is used for training. It also adopts the masking mechanism in action probability distribution for only selecting physical nodes with enough computing resources/ Particularly, during training, if encountering failure samples, customized negative rewards are returned. It also uses the probability distribution action masking mechanism. Similar to A3C-GCN, it introduces the penalty into reward shaping.
  • GAL-VNE (Geng et al., 2023) is a model-free RL method that formulates VNE into both Global learning across requests And Local prediction within requests (GAL). It first uses supervised learning to develop a GNN policy to solve VNE in one shot. Then, in online perspective, they use RL to finetune this policy to improve overall performance.
@@ F.3 Variations Descriptions
  • $\text{CONAL}_{\text{w/o ARB}}$ removes the proposed adaptive reachability budget (ARB) and uses zero as a fixed reachability budget.
@@ Appendix G Addtional Evaluation
In this section, we present additional experiments to evaluate the training stability, generalizability, scalability, and practicability of CONAL, as well as investigate the impact of two key hyperparameters.
G.1 Training Stability Study
To study the training stability of CONAL, we first provide an analysis of how different training conditions impact testing performance. Then, we dive into the convergence comparison between CONAL and $\text{CONAL}_{\text{w/o ARB}}$.
Figure 4: Results on testing performance ($\eta$ = 0.14) of algorithms trained under different average arrival rate $\eta$.
Figure 5: Training convergence curves of reward and Lagrange multiplier ($\eta$ = 0.26 for training).
Training Conditions vs. Testing Performance Analysis. Similar to our preliminary study in Appendix C, we further explore the impact of training conditions on CONAL performance. We train CONAL, $\text{CONAL}_{\text{w/o ARB}}$ and A3C-GCN under different arrival rates $\lambda$ of VN requests from 0.14 to 0.26, as higher arrival rates increase the frequency of unsolvable instances. After training, we evaluate the models using a fixed arrival rate of $\lambda=0.14$ with a random seed of 0 as the benchmark. The results are shown in Figures 4(b)(c)(d). We observe that as the arrival rate $\lambda$ of training conditions increases, CONAL maintains a more stable testing performance while $\text{CONAL}_{\text{w/o ARB}}$ and A3C-GCN shows an noticeable decline. Specifically, when trained under $\lambda=0.26$, which includes more unsolvable instances, two metrics (VN_RAC and LT_REV) of CONALw/o ARB even perform worse than A3C-GCN, which incorporates a penalty in the reward function. This may be because, while $\text{CONAL}_{\text{w/o ARB}}$ can perceive fine-grained constraints and aims to achieve zero violation, without the ARB module and the penalty in the reward function, it struggles to handle the increased number of failure samples caused by unsolvable instances effectively. This analysis shows that CONAL with the proposed ARB module is more stable in learning a high-quality policy by efficiently handling unsolvable instances and perceiving complex constraints.
Training Convergence Analysis. To investigate the training convergence of CONAL and $\text{CONAL}_{\text{w/o ARB}}$, particularly training with a high proportion of unsolvable instances, we conduct experiments with an arrival rate $\eta=0.26$ for VN requests on WX100, which corresponds to a scenario with more unsolvable instances. During training, we monitor both reward and Lagrange Multiplier over 300 training steps. (a) Learning Curve Comparison. We first compare the learning curves of CONAL and $\text{CONAL}_{\text{w/o ARB}}$, depicted in Figure 5(a). We observe that compared to $\text{CONAL}_{\text{w/o ARB}}$ fluctuates during the training process, CONAL exhibits greater stability and converges to better performance This stability and superior performance are attributed to our adaptive reachability budget method, which effectively addresses unsolvable instances during training, thus enhancing overall training stability and mitigating negative impact on performance. (b) Convergence Analysis of Lagrange Multiplier. As shown in Figure 5(b), the results reveal a clear distinction in the convergence behavior of $\lambda$. Without ARB, $\lambda$ diverges rapidly to extreme values. This divergence reflects instability in the optimization process and leads to unreliable policy updates. In contrast, CONAL with ARB effectively stabilizes $\lambda$ throughout training, close to a small value. By dynamically adjusting the feasibility budget, ARB mitigates the impact of unsolvable instances, ensuring stable training even under challenging conditions. These findings demonstrate the crucial role of ARB module in preventing numerical instability and enabling robust policy learning in scenarios with high unsolvability.
G.2 Generalizability Study
@@
Dynamic Request Distribution Testing. In practical scenarios, the node size and resource requirements distributions of VN requests may vary due to different service demands. To simulate such situations, similar to previous work (Geng et al., 2023), we equally divided 1000 VN requests into four sub-groups. In comparison to the default simulation settings, we modified one parameter related to the distribution of VN resource demand or VN node size for each sub-group: For the first group, the node and link resource distributions of VN are changed to [0, 30] and [0, 75], respectively; For the second group, the node and link resource distributions of VN are adjusted to [0, 40] and [0, 100], respectively; For the third group, the VN node size distribution is changed to [2, 15]; For the fourth group, the VN node size distribution is altered to [2, 20]. We evaluate the CONAL and other RL-based methods trained in the default settings. The results are illustrated in Figure 7. We observe that CONAL exhibits superior performance across all metrics, indicating its strong adaptability to dynamic request distributions. Regarding the VN_ACR, all algorithms experience a rapid decrease in the early stages of the simulation due to the quick consumption of resources. Subsequently, their VN_ACR becomes relatively stable. It is worth noting that in the fourth stage, most algorithms show a clear downward trend in VN_ACR. This is because embedding larger VNs with more complex topologies and greater resource demands becomes increasingly challenging. This testing shows that CONAL is highly adaptable across varying VN request distributions, which underscores CONAL’s generalization in dynamic systems.